Mathematics 243 – Algebraic Structures Definitions Quiz 5 – October 19, 2012

Your Name:

A) (5) If a, b ∈ Z, what does it mean to say d = gcd(a, b)? (Give our definition, not a method for computing d, or another characterization of d.)

B) (5) Is the divisibility relation on Z transitive? Why or why not?